# DWTNeRF: Boosting Few-shot Neural Radiance Fields via Discrete Wavelet Transform

### v2

@@ 1 Introduction
The above discussion yields a two-fold motivation. Firstly, given the proven effectiveness of frequency-based regularization, we need to devise a method that accomplishes the same idea, but compatible with INGP. To this end, we decompose the rendered and ground-truth scenes into levels of frequencies, using the Discrete Wavelet Transform. Their frequency-space differences are minimized by our novel Discrete Wavelet loss. By setting a higher weight to low-frequency discrepancies, we can explicitly prioritize their learning right in the training objective. Secondly, given the proven effectiveness of model-based methods, we need to devise a method that does not degrade INGP, which is sensitive to architectural changes. Our focus is on modelling the color-density interactions. While better modelled with separate MLP branches, they still require some levels of interactions [46]. We call this the “cross-branch interactions”. To this end, we propose using multi-head attention modules at the input and output levels. This does not incur architectural changes to the MLP, and does not degrade performance like other model methods.
@@
  • Regularization-based approach with a novel Discrete Wavelet loss that explicitly decomposes the scenes into low and high-frequencies, enabling prioritizing low frequencies at earlier training stages;
  • Model-based approach to enable cross-branch interactions using multi-head attention, which is compatible with fast-converging INGP;
@@ 2.2 Few-shot NeRF
Prior-based methods Those enable few-shot NeRF by training a generalized model on a large dataset of diverse scenes or by incorporating pre-trained priors in their training objectives. pixelNeRF [42] leveraged a feature encoder to extract image features from sparse input views, and conditioned a NeRF on these features. DietNeRF [11] introduced a semantic consistency loss that minimizes discrepancies in embeddings of rendered and ground-truth views, encouraging high-level semantic similarities. RegNeRF [23] leveraged a trained normalizing flow model to regularize color patches of unobserved views. SparseNeRF [8] introduced a local depth ranking loss based on priors from a trained depth estimation model. SPARF [33] minimized coordinate-space discrepancies between pixel matches in multiple ground-truth views, predicted by a trained feature matching model. This encourages a global and geometrically accurate solution. GeCoNeRF [17] extended this concept of geometric consistency to the feature-space, regularizing both semantically and structurally.
Regularization-based methods Those enable few-shot NeRF by additionally introducing regularization terms, optimizing in a per-scene manner. InfoNeRF [16] introduced a KL-divergence loss that enforces consistent density distributions across neighboring rays. RegNeRF [23] introduced a depth smoothness loss that penalizes depth discrepancies between neighboring points. Mip-NeRF 360 [1] introduced a distortion loss that reduces floating artifacts by minimizing the weighted distances between sampled points. DiffusioNeRF [37] introduced a full geometry loss that encourages the weights of sampled points along each ray to sum to unity, which ensures the rays are absorbed fully by the scene’s geometry. Another significant work is FreeNeRF [39], which did not introduce new terms but proposed masking input encodings in a coarse-to-fine manner, preventing NeRF from overfitting on high frequencies in early training stages. CombiNeRF [2], which our work is based on, synergistically combined all these techniques on top of INGP, achieving SOTA results.
Model-based methods [46] shows that Vanilla MLP is not sufficient for few-shot NeRF. They introduced two main modules, which we term “Residual Connections” and “Element-wise Cross-branch Interactions”. In the former, the encoded inputs are fed into each MLP layer, ensuring a shorter connection between the inputs and outputs. This is similar to ResNet [9] & DenseNet’s [10] residual connections. In the latter, the colors & densities are modeled using separate MLP branches, and cross-branch interactions are learned by a simple element-wise addition between corresponding layers. This pioneering work presented a new direction towards few-shot NeRF: adjusting Vanilla MLP’s architecture.
@@ 3.1 Neural Radiance Fields
$\mathcal{L}_{MSE}=\frac{1}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}||\mathbf{c}(\mathbf{r})-\mathbf{c}_{gt}(\mathbf{r})||_{2}^{2}$
@@ 3.2 Discrete Wavelet Transform (DWT)
where $\mathbf{L}$ and $\mathbf{H}$ are the low- and high-pass 2D matrices, respectively. $\mathbf{L}$ is constructed with shifting rows of its corresponding 1D low-pass wavelet filter $\bm{\ell}$, and similarly for $\mathbf{H}$. If $\bm{\ell}$ is orthogonal, then the 1D high-pass filter $\bm{h}$ is derived trivially from $\bm{\ell}$, and $\mathbf{H}$ trivially from $\mathbf{L}$ [30]. $\mathbf{I}^{LL}$ is called the LL sub-band of $\mathbf{I}$, and so on for other components. Intuitively, $\mathbf{I}^{LL}$ represents the low-frequency components of the image. $\mathbf{I}^{LH}$ and $\mathbf{I}^{HL}$ represent high frequencies in one direction, but low frequencies in another. Finally, $\mathbf{I}^{HH}$ represents the high frequencies in both vertical & horizontal directions. Because $\mathbf{L}$ is constructed with shifting rows of $\bm{\ell}$ (and similarly for $\mathbf{H}$), the sub-bands are only a quarter of the original image. This is a concern in applying the DW loss, which we will explain later.
@@ 4.1 Overview
Figure 2 shows the architecture diagram of our DWTNeRF. It consists of two main modules: the Discrete Wavelet (DW) loss and cross-branch interactions. The DW loss decomposes rendered and ground-truth views into levels of frequencies, and minimizes the differences in their frequency representations. By setting a higher weight for low-frequency differences, we can explicitly prioritize them in earlier training stages, thereby reducing overfitting on high frequencies. This represents a regularization-based approach that is frequency-centric. On the other hand, the cross-branch modules utilize multi-head attention to capture interactions not only between colors and densities but also among neighboring 3D points. This helps mitigate overfitting to high-frequency details by constraining color predictions through density-based, position-aware relationships. This also represents a model-based method that is compatible with INGP, which is highly sensitive to architectural changes.
@@ 4.2 Scene Decomposition with Discrete Wavelet loss
where $\mathcal{L}_{dist}$ is the distortion loss, introduced by Mip-NeRF 360 [1]. $\mathcal{L}_{fg}$ is the full geometry loss, introduced by DiffusioNeRF [37]. $\mathcal{L}_{ds}$ is the depth smoothness loss, introduced by RegNeRF [23]. $\mathcal{L}_{KL}$ is the KL divergence loss, introduced by InfoNeRF [16]. The DW loss allows us to prioritize low frequencies directly in the training objective.
@@
Patch-based Rendering The DW loss requires adjacently sampled rays $\mathbf{r}$ that make up a 2D image $\mathbf{I}$. This requires patch-based ray sampling. As illustrated in Figure 3b, rays sampled in patches are adjacent to each other, making up a small square grid. Figure 3a shows rays sampled randomly throughout the image grid. We use random ray sampling for calculating $\mathcal{L}_{MSE}$, and patch-based for $\mathcal{L}_{DW}$. This design choice is intentional: it considers both global structures (random rays) and local structures (patch-based rays) in one iteration.
Lazy Regularization The sub-bands are 4 times smaller than the original image $\mathbf{I}$. For the DW loss to be feasible, however, the sub-bands still need to be visually meaningful. For example, $\mathbf{I}^{LH}$ and $\mathbf{I}^{HL}$ can highlight vertical and horizontal edges of a scene. By minimizing their differences with the ground-truths, we can better preserve those mid-frequency-level structures. This requires the rendered image $\mathbf{I}$ to be large in the first place, or else its sub-bands would be too coarse. This incurs heavy computational overhead, because an $N$-sized square patch requires $N^{2}$ number of rays. We alleviate the problem as follows. Firstly, we follow the lazy regularization procedure inspired by [14], where we only calculates the DW loss every few iterations. Secondly, since we only prioritize low frequencies earlier in the training, we can safely stop calculating the DW loss later on. We present the full training algorithm of DWTNeRF at Algorithm 1. Here, we only detail $\mathcal{L}_{MSE}$ and $\mathcal{L}_{DW}$ for simplicity.
@@ 4.3 Cross-branch Interactions with Multi-head Attention
Our approach for cross-branch interactions consists of:
@@
I. Cross-branch Concatenation We are given the hash encoding $\gamma_{0:N_{\mathbf{x}}}(\mathbf{x})$ and the view encoding $\gamma_{0:N_{\mathbf{d}}}(\mathbf{d})$. Here, the subscripts denote the dimension of the encoding. $0:N_{(\cdot)}$ means the full dimension. $N_{\mathbf{x}}$ and $N_{\mathbf{d}}$ represent the highest dimensions of the hash and view encodings. The first stage involves concatenating the view encoding with all-but-the-first-dimension of the hash encoding. This can be written as $\gamma^{\prime}(\mathbf{d},\mathbf{x})=\left[\gamma(\mathbf{d}),\gamma_{1:N_{\mathbf{x}}}(\mathbf{x})\right]$, where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operator. We call this the “hybrid encoding”. The hash encoding only retains its first dimension: $\gamma^{\prime}(\mathbf{x})=\gamma_{0}(\mathbf{x})$. The intuition is as follows. We will use $\gamma^{\prime}(\mathbf{x})$ as inputs to the density MLP, and $\gamma^{\prime}(\mathbf{d},\mathbf{x})$ as inputs to the color MLP. Densities are irrespective of viewing directions and should only be position-dependent, so we make $\gamma^{\prime}(\mathbf{x})$ extremely lean to avoid over-parameterizing them. On the other hand, colors are both view- and position-dependent, which explains the concatenation.
@@
II. Input-level Multi-head Attention We now have the “trimmed” hash encoding $\gamma^{\prime}(\mathbf{x})$ and hybrid encoding $\gamma^{\prime}(\mathbf{d},\mathbf{x})$. We push $\gamma^{\prime}(\mathbf{d},\mathbf{x})$ into a multi-head attention module to learn the interactions between $\gamma(\mathbf{d})$ and $\gamma_{1:N_{\mathbf{x}}}(\mathbf{x})$. Again, we leave $\gamma^{\prime}(\mathbf{x})$ intact to not over-parameterize it. The multi-head attention is applied as: $\texttt{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})=[\texttt{head}_{0},...,\texttt{head}_{H}]\mathbf{W}^{O}$, where $\texttt{head}_{i}=\texttt{Attn}(\mathbf{Q}\mathbf{W}_{i}^{\mathbf{Q}},\mathbf{K}\mathbf{W}_{i}^{\mathbf{K}},\mathbf{V}\mathbf{W}_{i}^{\mathbf{V}})$. Here, Attn is simply the attention layer [34]. $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ are the query, key and value matrices, all of which are set to $\gamma^{\prime}(\mathbf{d},\mathbf{x})$. $\mathbf{W}_{i}^{\mathbf{Q}}$, $\mathbf{W}_{i}^{\mathbf{K}}$, $\mathbf{W}_{i}^{\mathbf{V}}$ are the corresponding learnable projection matrices. We note that $\gamma^{\prime}(\mathbf{d},\mathbf{x})$ is a tensor of size ||×$|\mathcal{R}|\times(N_{\mathbf{x}}-1+N_{\mathbf{d}})$. The choice of using attention is deliberate. Along the vertical dimension, it learns the interactions between all rays $\mathbf{r}$ in the batch $\mathcal{R}$. This equates to learning the interactions between neighboring 3D points. Along the horizontal dimension, it learns the interactions between the positions and viewing directions. Using many heads, up to a maximum $H$, provides many of such interactions. The heads are concatenated, and then projected by another learnable matrix $\mathbf{W}^{O}$ to ensure outputs and inputs are of the same size. We now write the attended hybrid encoding as $\gamma^{\prime\prime}(\mathbf{d},\mathbf{x})$. The plenoptic function is now $f(\gamma^{\prime}(\mathbf{x}),\gamma^{\prime\prime}(\mathbf{d},\mathbf{x}))=\sigma,\mathbf{c}$, parameterized by two separate MLP branches for densities and colors.
@@
III. Output-level Multi-head Attention $f$’s outputs are the “preliminary” colors $\mathbf{c}$ and densities $\sigma$. We again learn their interactions. We concatenate them together, and feed them into a multi-head attention module. Here, $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ are set to $\bm{\zeta}=[\sigma,\mathbf{c}]$. All intuitions in the second stage still hold. We “un-concatenate” the attended $\bm{\zeta}^{\prime}$ to obtain the final predictions. Specifically, the densities $\sigma^{\prime}$ are the first dimension of $\bm{\zeta}^{\prime}$, and the colors $\mathbf{c}^{\prime}$ are the three remaining dimensions.
@@ 5.1 Datasets & Implementation Details
Datasets & Metrics DWTNeRF is evaluated on the LLFF [20] dataset and the NeRF-Synthetic dataset [21], under few-shot conditions. Those are popular benchmarks for novel view synthesis. For the LLFF dataset, we evaluated the test metrics under 2/3/6/9 input views. For the NeRF-Synthetic (NS) dataset, we trained on 4 views and tested on 25 views. Those protocols are similar to previous works. We employed the PSNR, SSIM [35] and LPIPS [43] metrics to evaluate the quality of novel views.
@@
Implementation Details DWTNeRF is built on top of CombiNeRF [2], which in turn is built on a Pytorch implementation of INGP, torch-ngp [31]. We kept all default settings of CombiNeRF, and trained for 10K iterations in all experiments. We randomly sampled 4096 rays (for LLFF) and 7008 rays (for NS) in each iteration. For DW loss, we additionally sampled 36864 rays, leading to 192 $\times$ 192 square patches. We calculated the DW loss every 10 (for LLFF) and 150 (for NS) iterations, but stopped altogether at 5K iterations to facilitate high-frequency learning. For LLFF, we set the LL sub-band weight as 0.4, the other sub-bands were 0.2. For NS, LL sub-band weight was 0.04, the rest were all 0.02. This ensures that the low frequencies were the most important. Finally, we used 2 (for LLFF) and 1 (for NS) attention heads for both input and output levels. All trainings were done with GeForce RTX 4070 Ti Super.

### v3

@@ 1 Introduction
The above discussion yields a two-fold motivation. Firstly, given the proven effectiveness of frequency-based regularization, we need to devise a method that accomplishes the same idea, but compatible with INGP. To this end, we decompose the rendered and ground-truth scenes into levels of frequencies, using the Discrete Wavelet Transform. Their frequency-space differences are minimized by our novel Discrete Wavelet loss. By setting a higher weight to low-frequency discrepancies, we can explicitly emphasize their learning right in the training objective. Secondly, given the proven effectiveness of model-based methods, we need to devise a method that does not degrade INGP, which is sensitive to architectural changes. Our focus is on modelling the color-density interactions. While better modelled with separate MLP branches, they still require some levels of interactions [46]. We call this the “cross-branch interactions”. To this end, we propose using multi-head attention modules at the input and output levels. This does not incur architectural changes to the MLP, and does not degrade performance like other model-based methods.
@@
  • Regularization-based approach with a novel Discrete Wavelet loss that explicitly decomposes the scenes into low and high-frequencies, emphasizing low frequencies at earlier training stages.
  • Model-based approach to enable color-density interactions using multi-head attention, which is compatible with fast-converging INGP;
@@ 2.2 Few-shot NeRF
Prior-based methods. Those enable few-shot NeRF by training a generalized model on a large dataset of diverse scenes or by incorporating pre-trained priors in their training objectives. pixelNeRF [42] leveraged a feature encoder to extract image features from sparse input views, and conditioned a NeRF on these features. DietNeRF [11] introduced a semantic consistency loss that minimizes discrepancies in embeddings of rendered and ground-truth views, encouraging high-level semantic similarities. RegNeRF [23] leveraged a trained normalizing flow model to regularize color patches of unobserved views. SparseNeRF [8] introduced a local depth ranking loss based on priors from a trained depth estimation model. SPARF [33] minimized coordinate-space discrepancies between pixel matches in multiple ground-truth views, predicted by a trained feature matching model. This encourages a global and geometrically accurate solution. GeCoNeRF [17] extended this concept of geometric consistency to the feature-space, regularizing both semantically and structurally.
Regularization-based methods. Those enable few-shot NeRF by additionally introducing regularization terms, optimizing in a per-scene manner. InfoNeRF [16] introduced a KL-divergence loss that enforces consistent density distributions across neighboring rays. RegNeRF [23] introduced a depth smoothness loss that penalizes depth discrepancies between neighboring points. Mip-NeRF 360 [1] introduced a distortion loss that reduces floating artifacts by minimizing the weighted distances between sampled points. DiffusioNeRF [37] introduced a full geometry loss that encourages the weights of sampled points along each ray to sum to unity, which ensures the rays are absorbed fully by the scene’s geometry. Another significant work is FreeNeRF [39], which did not introduce new terms but proposed masking input encodings in a coarse-to-fine manner, preventing NeRF from overfitting on high frequencies in early training stages. CombiNeRF [2], which our work is based on, synergistically combined all these techniques on top of INGP, achieving SOTA results.
Model-based methods. [46] shows that Vanilla MLP is not sufficient for few-shot NeRF. They introduced two main modules, which we term “Residual Connections” and “Element-wise Cross-branch Interactions”. In the former, the encoded inputs are fed into each MLP layer, ensuring a shorter connection between the inputs and outputs. This is similar to ResNet [9] & DenseNet’s [10] residual connections. In the latter, the colors & densities are modeled using separate MLP branches, and cross-branch interactions are learned by a simple element-wise addition between corresponding layers. This pioneering work presented a new direction towards few-shot NeRF: adjusting Vanilla MLP’s architecture.
@@ 3.1 Neural Radiance Fields
$\mathbf{c}(\mathbf{r})=\int_{t_{n}}^{t_{f}}T(t)\sigma(\mathbf{r}(t))\mathbf{c}(\mathbf{r}(t),\mathbf{d})\,dt\$(1)
@@ 3.2 Discrete Wavelet Transform (DWT)
[…]
@@ 4.1 Overview
Figure 2 shows the architecture diagram of our DWTNeRF. It consists of two main modules: the Discrete Wavelet (DW) loss and cross-branch interactions. The DW loss decomposes rendered and ground-truth views into levels of frequencies, and minimizes the differences in their frequency representations. By setting a higher weight for low-frequency differences, we can explicitly emphasize them in earlier training stages, thereby reducing overfitting on high frequencies. This represents a regularization-based approach that is frequency-centric. On the other hand, the cross-branch modules utilize multi-head attention to capture interactions not only between colors and densities but also among neighboring 3D points. This helps mitigate overfitting to high-frequency details by constraining color predictions through density-based, position-aware relationships. This also represents a model-based method that is compatible with INGP, which is highly sensitive to architectural changes.
@@ 4.2 Scene Decomposition with Discrete Wavelet loss
where $\mathcal{L}_{dist}$ is the distortion loss, introduced by Mip-NeRF 360 [1]. $\mathcal{L}_{fg}$ is the full geometry loss, introduced by DiffusioNeRF [37]. $\mathcal{L}_{ds}$ is the depth smoothness loss, introduced by RegNeRF [23]. $\mathcal{L}_{KL}$ is the KL divergence loss, introduced by InfoNeRF [16]. The DW loss allows us to emphasize low frequencies directly in the training objective.
@@
Patch-based Rendering. The DW loss requires adjacently sampled rays $\mathbf{r}$ that make up a 2D image $\mathbf{I}$. This requires patch-based ray sampling. As illustrated in Figure 3b, rays sampled in patches are adjacent to each other, making up a small square grid. Figure 3a shows rays sampled randomly throughout the image grid. We use random ray sampling for calculating $\mathcal{L}_{MSE}$, and patch-based for $\mathcal{L}_{DW}$. This design choice is intentional: it considers both global structures (random rays) and local structures (patch-based rays) in one iteration.
Lazy Regularization. The sub-bands are 4 times smaller than the original image $\mathbf{I}$. For the DW loss to be feasible, however, the sub-bands still need to be visually meaningful. For example, $\mathbf{I}^{LH}$ and $\mathbf{I}^{HL}$ can highlight vertical and horizontal edges of a scene. By minimizing their differences with the ground-truths, we can better preserve those mid-frequency-level structures. This requires the rendered image $\mathbf{I}$ to be large in the first place, or else its sub-bands would be too coarse. This incurs heavy computational overhead, because an $N$-sized square patch requires $N^{2}$ number of rays. We alleviate the problem as follows. Firstly, we follow the lazy regularization procedure inspired by [14], where we only calculates the DW loss every few iterations. Secondly, since we only prioritize low frequencies earlier in the training, we can safely stop calculating the DW loss later on. We present the full training algorithm of DWTNeRF at Algorithm 1. Here, we only detail $\mathcal{L}_{MSE}$ and $\mathcal{L}_{DW}$ for simplicity.
@@ 4.3 Cross-branch Interactions with Multi-head Attention
Our approach for cross-branch interactions consists of the following:
@@
I. Cross-branch Concatenation. We are given the hash encoding $\gamma_{0:N_{\mathbf{x}}}(\mathbf{x})$ and the view encoding $\gamma_{0:N_{\mathbf{d}}}(\mathbf{d})$. Here, the subscripts denote the dimension of the encoding. $0:N_{(\cdot)}$ means the full dimension. $N_{\mathbf{x}}$ and $N_{\mathbf{d}}$ represent the highest dimensions of the hash and view encodings. The first stage involves concatenating the view encoding with all-but-the-first-dimension of the hash encoding. This can be written as $\gamma^{\prime}(\mathbf{d},\mathbf{x})=\left[\gamma(\mathbf{d}),\gamma_{1:N_{\mathbf{x}}}(\mathbf{x})\right]$, where $\cdot$$\cdot$] denotes the concatenation operator. We call this the “hybrid encoding”. The hash encoding only retains its first dimension: $\gamma^{\prime}(\mathbf{x})=\gamma_{0}(\mathbf{x})$. The intuition is as follows. We will use $\gamma^{\prime}(\mathbf{x})$ as inputs to the density MLP, and $\gamma^{\prime}(\mathbf{d},\mathbf{x})$ as inputs to the color MLP. Densities are irrespective of viewing directions and should only be position-dependent, so we make $\gamma^{\prime}(\mathbf{x})$ extremely lean to avoid over-parameterizing them. On the other hand, colors are both view- and position-dependent, which explains the concatenation.
@@
II. Input-level Multi-head Attention. We now have the “trimmed” hash encoding $\gamma^{\prime}(\mathbf{x})$ and hybrid encoding $\gamma^{\prime}(\mathbf{d},\mathbf{x})$. We push $\gamma^{\prime}(\mathbf{d},\mathbf{x})$ into a multi-head attention module to learn the interactions between $\gamma(\mathbf{d})$ and $\gamma_{1:N_{\mathbf{x}}}(\mathbf{x})$. Again, we leave $\gamma^{\prime}(\mathbf{x})$ intact to not over-parameterize it. The multi-head attention is applied as: $\texttt{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})=[\texttt{head}_{0},...,\texttt{head}_{H}]\mathbf{W}^{O}$, where $\texttt{head}_{i}=\texttt{Attn}(\mathbf{Q}\mathbf{W}_{i}^{\mathbf{Q}},\mathbf{K}\mathbf{W}_{i}^{\mathbf{K}},\mathbf{V}\mathbf{W}_{i}^{\mathbf{V}})$. Here, Attn is simply the attention layer [34]. $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ are the query, key and value matrices, all of which are set to $\gamma^{\prime}(\mathbf{d},\mathbf{x})$. $\mathbf{W}_{i}^{\mathbf{Q}}$, $\mathbf{W}_{i}^{\mathbf{K}}$, $\mathbf{W}_{i}^{\mathbf{V}}$ are the corresponding learnable projection matrices. We note that $\gamma^{\prime}(\mathbf{d},\mathbf{x})$ is a tensor of size $|\mathcal{R}|\times(N_{\mathbf{x}}-1+N_{\mathbf{d}})$. The choice of using attention is deliberate. Along the vertical dimension, it learns the interactions between all rays $\mathbf{r}$ in the batch $\mathcal{R}$. This equates to learning the interactions between neighboring 3D points. Along the horizontal dimension, it learns the interactions between the positions and viewing directions. Using many heads, up to a maximum $H$, provides many of such interactions. The heads are concatenated, and then projected by another learnable matrix $\mathbf{W}^{O}$ to ensure outputs and inputs are of the same size. We now write the attended hybrid encoding as $\gamma^{\prime\prime}(\mathbf{d},\mathbf{x})$. The plenoptic function is now $f(\gamma^{\prime}(\mathbf{x}),\gamma^{\prime\prime}(\mathbf{d},\mathbf{x}))=\sigma,\mathbf{c}$, parameterized by two separate MLP branches for densities and colors.
@@
III. Output-level Multi-head Attention. $f$’s outputs are the “preliminary” colors $\mathbf{c}$ and densities $\sigma$. We again learn their interactions. We concatenate them together, and feed them into a multi-head attention module. Here, $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ are set to $\bm{\zeta}=[\sigma,\mathbf{c}]$. All intuitions in the second stage still hold. We “un-concatenate” the attended $\bm{\zeta}^{\prime}$ to obtain the final predictions. Specifically, the densities $\sigma^{\prime}$ are the first dimension of $\bm{\zeta}^{\prime}$, and the colors $\mathbf{c}^{\prime}$ are the three remaining dimensions.
@@ 5.1 Datasets & Implementation Details
Datasets & Metrics. DWTNeRF is evaluated on the LLFF [20] dataset and the NeRF-Synthetic dataset [21], under few-shot conditions. Those are popular benchmarks for novel view synthesis. For the LLFF dataset, we evaluated the test metrics under 2/3/6/9 input views. For the NeRF-Synthetic (NS) dataset, we trained on 4 views and tested on 25 views. Those protocols are similar to previous works. We employed the PSNR, SSIM [35] and LPIPS [43] metrics to evaluate the quality of novel views.
@@
Implementation Details. DWTNeRF is built on top of CombiNeRF [2], which in turn is built on a Pytorch implementation of INGP, torch-ngp [31]. We kept all default settings of CombiNeRF, and trained for 10K iterations in all experiments. We randomly sampled 4096 rays (for LLFF) and 7008 rays (for NS) in each iteration. For DW loss, we additionally sampled 36864 rays, leading to 192 $\times$ 192 square patches. We calculated the DW loss every 10 (for LLFF) and 150 (for NS) iterations, but stopped altogether at 5K iterations to facilitate high-frequency learning. For LLFF, we set the LL sub-band weight as 0.4, the other sub-bands were 0.2. For NS, LL sub-band weight was 0.04, the rest were all 0.02. This ensures that the low frequencies were the most important. Finally, we used 2 (for LLFF) and 1 (for NS) attention heads for both input and output levels. All trainings were done with GeForce RTX 4070 Ti Super.